LCG Grid Deployment - gLite Pre Production Services - CERN PPS site

CERN PPS Work Log

CERN_PPS (C5 Report on  2 April-09)
CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
Site upgraded to latest update of gLite 3.2.0 PPS update 01.


CERN_PPS (C5 Report on  12th March-09)

CERN GRID Pre-Production Site (CERN_PPS):  
-----------------------------------------
* gLite3.1.0-PPS-UPDATE45 was deployed. All services were upgraded successfully.
* Latest version of lcg-CAs v1.28-1 tested.
* Installation of glite-WN (SLC5 64bit) on progress.


CERN_PPS (C5 Report on  19th-Feb-09)
CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
1. gLite 3.1 PPS Update 44 was released this week and deployed on:
- glite-WN
- lcg-CE
- glite-AMGA (oracle)
- CREAM
- DPM Mysql
Minor issue in yaim configuration for glite-UI was found. Bug#40653 reopened.

2. PPS AFS UI upgraded to latest version v-3.1.30
3. EGEE-SEE-CERT has started to publish in PPS. Because the site wanted to be a part of PPS to make certification more faster
and to get more traffic to there resources.
4. There is no SAM submission to PPS since last week. We are trying to contact responsible unit.

  • 2009-02-17 : Task: EGEE-SEE-CERT is added to CERN_PPS BDII. Details:Main reason was, site wanted to be part of PPS to make certification more faster and to get more traffic to their CREAM CE. Status:Done Farida

  • 2009-02-17 : Task: Pre-deployment test done for glite-WN and lcg-CE and AMGA-Server. Details:Found configuration issue in glite-WN. Status:Done Farida

  • 2009-02-18 : Task: PPS-AFS UI updated to latest version. Details: An issue in yaim Configuration. Patch#40653 reopened. Workaround applied to UI.Status:Done Farida


CERN_PPS (C5 Report on  5th-Feb-09)

CERN GRID Pre-Production Site (CERN_PPS):  
-----------------------------------------
* PPS AFS UI upgraded to the latest version 3.1.29
* Latest version of lcg-CAs v1.27-1 tested.


CERN_PPS (C5 Report on  29-Jan-09)

CERN GRID Pre-Production Site (CERN_PPS):  
-----------------------------------------
* gLite3.1.0-PPS-UPDATE43, was released yesterday and deployed on:
- FTS/FTA
- top-level BDII
-glite-WN
-CREAM CE
A small  yaim configuration issue was found. See BUG:46400 

* Problem in publishing pps-wms. After digging, latest version of info provider make WMS and list of supported VOs disappear from information system. 
  For details BUG:46209
* LSF CE debugging. For some unknown reasons log rotation didn't clean up lsflogs and host ended up with No space left on device. However manual fix has been applied.
* Kernel upgraded to 2.6.9-78.0.13 to preproduction cluster.


CERN_PPS (C5 Report on  15-Jan-09)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* pre-deployment test of gLite 3.1 PPS Update 42 Done:
- glite-AMGA (oracle)
- CREAM
- glite-WN
- DPM Mysql

* PPS AFS UI upgraded to latest version 3.1.28
* CREAM pilot resources of INFN-CNAF and INFN-BARI published by CERN_PPS site-BDII inorder to test them by SAM..
* Malfunctioning of LB server affecting job submission via WMS pps-wms.cern.ch after PPS update41.
  - Bugs already reported.  
  - To fix bugs, a patch https://savannah.cern.ch/patch/index.php?2562 is incertification with High-priority.
  - WE are in touch with developers to get a workaround to fix the problem. 
   

CERN_PPS (C5 Report on  4th-Dec-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
 gLite3.1.0 PPS Update 41 has been done on lcg-CE, site-BDII.


CERN_PPS (C5 Report on  30-October-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
pre-deployment test of gLite 3.1 PPS Update 38
- FTS/FTA: Done
- CREAM: Done
- glite-WN
- DPM Mysql
An issue regarding  messing up of voms dir in yaim function of CREAM was found.

PPS AFS UI upgraded to the latest version 3.1.23



CERN_PPS (C5 Report on  15-October-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
   * The BDII run on behalf of the Cern ROC was updated to glite 3.1 Update33
      - An issue was found reported with BUG:42799 (see PPS report)
   * The FTS/FTA is set-up and it is now fully functional at CERN_PPS
   * The PPS SAM server was upgraded to the last release
 



CERN_PPS (C5 Report on  01-October-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
pre-deployment test of gLite 3.1 PPS Update 37 
- FTS/FTA: Done
- CREAM: Done
- BDII: Done

test of last CA Update performed

  



CERN_PPS (C5 Report on  18-September-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
pre-deployment test of gLite 3.1 PPS Update 36 in progress
  



CERN_PPS (C5 Report on  4-September-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

nothing to report
  



CERN_PPS (C5 Report on  28 August-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

nothing to report
  


--

CERN_PPS (C5 Report on  21 August-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

nothing to report
  


CERN_PPS (C5 Report on  14 August-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

* Pre-deployment test for PPS-update 35
   - FTS  and AMGA SERVER : DONE

* PPS AFS UI upgraded to gLite3.1.0 Update 34 
An issue  was found in YAIM configuration. Bug  was opened.
https://savannah.cern.ch/bugs/?func=detailitem&item_id=39910

  

--

CERN_PPS (C5 Report on  07 August-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

* Pre-deployment test for PPS-update 34
   - CREAM CE upgraded

* Version v1.24-1 of CAs tested and deployed

  


  • 2008-08-06 : Task: PPS-update 34 deployed on cream CE. Details: Upgrade and tests were fine.Status:Done Farida

  • 2008-08-01 : Task: CA (v1.24-1) updated and tested. Details: Deployed on CEs, DPM,glite-WNs Status:Done Farida


CERN_PPS (C5 Report on  31 July-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

* Pre-deployment test for PPS-update 33
   - on CREAM CE in progress

* PPS AFS UI upgraded to gLite3.1.0 Update 32

  


CERN_PPS (C5 Report on  24 July-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

* gLite3.1.0 PPS Update 32 pre-deployment test: 
   - pre-deployment test of glite-FTS and glite-FTA (Oracle) done.

* Installation of SAM server in progress
   - DB: done
   - webservice: done
   - bdii2oracle: done
   - portal: in progress

* Migration of site to new hardware in progress 


  • 2008-07-24 : Task: Migrating pps-ftm to a new hardware (lxbra2507) Details: Status:In progress Danica

  • 2008-07-22 : Task: New release of FTS_oracle and FTA_oracle for SL4 32 bit is installed on new hardware, and pre-deployment testing was done. Details: PPS-glite-FTS_oracle and PPS-glite-FTA_oracle (gLite3.1 PPS-Update32) installed and configured on lxbra2507.cern.ch. It replaced old pps-fts (lxb2064). Status:Done Danica


CERN_PPS (C5 Report on  10 July-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

* gLite3.1.0 PPS Update 31 pre-deployment test: 
   - glite-WN Done.

* Upgrade of the site to glite 3.1 PPS-update31 done.

* Installation of AMGA server on new hardware in progress.


  • 2008-07-04 : Task: General purpos node on lxb1908 migrated on lxbra2110 Details: Status:Done Antonio

CERN_PPS (C5 Report on  03 July-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------

* PPS AFS UI upgraded to gLite3.1.0 Update 27

* Re-installation of all the site on new hardware in progress
  Torque-server: done
  glite-WNs:  done
  lcg-CE (PBS): done
  DPM(MySQL): done
  top-level BDII: done
  Cream-CE: done


  • 2008-06-31 : Task: PPS AFS UI updated to pps update 25 Status:Done Farida

  • 2008-06-31 : Task: DPM MySQL installed on new hardware Details: Status:Done Farida

  • 2008-06-31 : Task: Install PPS Cluster on new hrdware Details: Status:Done Farida

  • 2008-06-31 : Task: Top level BDII installed on new machine Details: PPS-glite-BDII (gLite3.1 PPS-Update29) installed and configured as top level BDII on machine lxbra2506.cern.ch. It is ready to replace old pps-bdii (lxb2055). Status:Done Danica


CERN_PPS (C5 Report on  26-Jun-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* re-installation of all the site on new hardware in progress
* CERN_PPS is now operating a new instance of teh SAM server DB for PPS



CERN_PPS (C5 Report on  19-Jun-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* nothing to report.
* waiting for new hardware to replace servers out of warranty



CERN_PPS (C5 Report on  12-Jun-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to report



CERN_PPS (C5 Report on  5-Jun-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to report


CERN_PPS (C5 Report on 29-May-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Pre-deployment test for PPS-update 29
   - on top-BDII and site-BDII : done

* Pre-production instance of the AMGA server has been started.
 LHCb start running tests.

* pps-wms.cern.ch was reinstalled with SL4 and the latest WMS 
  for gLite 3.1 on SL4. Tested. Required workarounds applied.
  The services is OK now.

  • 2008-05-28 : Task: pps-wms.cern.ch was reinstalled with SL4 and latest WMS for gLite 3.1 Details: pps-wms.cern.ch was reinstalled with SL4 and the latest WMS for gLite 3.1 on SL4. Tested. Required workarounds applied. The services is OK now. Status:Done Konstantin

  • 2008-05-28 : Task: site-BDII upgraded to update29. Details: Test was OK.Status:Done Farida

  • 2008-05-28 : Task: pre-deployment test Details: gLite3.1 PPS-Update29: test of BDII upgrade successful. Report sent Status:Done Danica

  • 2008-05-26:Task: AMGA service for LHCb installation completed. Details: Preliminary Test was successful. LHCB has now started to run tests. Status:Done Farida


CERN_PPS (C5 Report on 22-May-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Site upgrading to gLite3.1.0 PPS Update 27
* Upgrade to gLite3.0.2 PPS Update 49 Completed



  • 2008-05-20: Task: WMS to PPS U49 Details: no problem Status: Done. Konstantin

  • 2008-05-15 : Task: pre-deployment test Details: gLite3.1 PPS-Update27: test of BDII upgrade successful. Report sent Status:Done Danica


CERN_PPS (C5 Report on 15-May-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Set-up of AMGA service for LHCb in progress
   - critical bug opened on configuration
     BUG:36599
     under investigation


  • 2008-05-15 : Task: AMGA for LHCb Details:critical BUG:36599 on configuration opened Status:In Progress Farida


CERN_PPS (C5 Report on 8-May-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Set-up of AMGA service for LHCb in progress




CERN_PPS (C5 Report on 24-April-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to report



CERN_PPS (C5 Report on 17-April-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update22
   top-BDII : Done

* Installation of AMGA server for LHCb in progress


  • 2008-04-15 : Task: AMGA for LHCb Details:lxn1187 dedicated to this installation, waiting for db parameters from Elisa. Status:In Progress Antonio

  • 2008-04-11 : Task: PPS-update 22 Details:top-BDII updated.Status:Done. Antonio


CERN_PPS (C5 Report on 10-April-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update22
   CEs, glite-WNs, site-BDII : Done

* PPS AFS UI upgraded to PPS-update22
  A security issue in YAIM configuration was found that block the running of configuration of the service. A Bug#35307 was already opened. However the manual fix was applied.
https://savannah.cern.ch/bugs/?35307



  • 2008-04-09 : Task: PPS-update 22 Details:CEs, WNs, site-BDII updated.Status:Done. Farida

  • 2008-03-20 : Task: PPS AFS UI update 22 Details: Security issue was found that block the configuration durring the update. see bug#35307. Manual fix was applied to latest patch of yaim. Status:Done. Farida


CERN_PPS (C5 Report on 3-Apr-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* A discussion was started with LHCb to set-up
  at CERN a pre-production instance of an AMGA server 
  for the VO based on the freshly released gLite distribution


  • 2008-03-20 : Task: ROC BDII Details:The system is now pointing to a static list in BDII_HTTP_URL="http://www.cern.ch/roc/bdii/static-bdii-cnaf-change-on-may-2008.conf", to be changed after CNAF down time. Status:Done. Antonio


CERN_PPS (C5 Report on 20-Mar-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update21
Done
* Latest version of CAs v1.20-1 tested and deployed
* OSG site started OSG-->EGEE interoperability tests 
  using resources in CERN_PPS
* The top-level BDII roc-bdii was re-installed and it is now 
   used as back-up for various operation applications 
   during CNAf downtime


  • 2008-03-20 : Task: ROC BDII Details:System upgraded to SL4. New version of BDII installed Status:Done. Antonio

  • 2008-03-19 : Task: PPS update 21 Details:Status:Done. Farida

  • 2008-03-19 : Task: latest version lcg-CA update 20 Details:testedStatus:Done. Farida


CERN_PPS (C5 Report on 13-Mar-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* AFS UI now using PPS-update19
* Accounting data is now published correctly for 
  CERN_PPS. A recipe on how to do it in a site with 
  multiple CEs is now available in 
  https://twiki.cern.ch/twiki/bin/view/LCG/PublishAccountingData

  • 2008-03-10 : Task: Accunting data publisher Details:Site is now publishing accounting data.Status:Done. Farida

  • 2008-03-10 : Task:PPS AFS UI update 19 Details: Status:Done. Farida


CERN_PPS (C5 Report on 6-Mar-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update19
Done

* Upgraded to gLite 3.0 PPS Update 46
  - WMS, FTS, MON box :done

CERN_PPS (C5 Report on 6-Mar-08)
---
CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to report.
   

CERN_PPS (C5 Report on 28-Feb-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Pre-deployment test for PPS-update 19
   - on LCF CE : done
* Accounting data publish for CERN_PPS in progress
* Job Priority test was done on LSF CE by CMS and ATLAS successful
* Planned to deploy glite-WN on 64 bit machines. Under discussion.
   

CERN_PPS (C5 Report on 14-Feb-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update18
   - CEs : done
   - DPM: done
   - glite-WNs: done
   - top BDII: done
   - site-BDII:done
   -  AFS UI for PPS : done

* Upgraded to gLite 3.0 PPS Update 45
  - WMS :done

* Support to CMS for test of Job Priority on LSF CE (in progress)
   

* 2008-02-19 : Task:PPS-update 18 Details: CEs, WNs, BDII, DPM updated. Status:Done. Farida

  • 2008-02-20: Task: AFS_UI for PPS update 18 Details: tested Status: Done. Farida
  • 2008-02-15: Task: WMS to PPS U45 Details: no problem Status: Done. Konstantin


CERN_PPS (C5 Report on 14-Feb-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* gLite3.1.0 PPS Update18 pre-deployment test: 
   - top-level BDII Done.
   
* FTS is upgraded to PPS-Update 44
* PPS AFS UI upgraded to gLite3.1.0 PPS Update17. Apparently no problem.
* CEs and WNs upgraded to gLite3.1.0 PPS Update17.


  • 2008-02-14 : Task:gLite3.1 U18 for pps top BDII Details:tested. no problem Status:Done. Lanxin

  • 2008-02-13 : Task:AFS UI for PPS updated 17 Details:tested. Apparently no problem Status:Done. Farida

  • 2008-02-10 : Task:PPS-update 17 Details:CEs, glite-WN updated Status:Done. Farida

CERN_PPS (C5 Report on 7-Feb-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* gLite3.1.0 PPS Update 17 pre-deployment test: 
   - top-level BDII Done.
   
* FTS is upgraded to PPS-Update 44
* PPS AFS UI upgraded to gLite3.1.6-1 (update 16). Broadcast send.
* Work in progress on publish accounting data for CERN_PPS site.


  • 2008-02-04 : Task:pps BDII to gLite 3.1 U17 Details: No problem Status:Done. Lanxin
  • 2008-02-01: Task: FTS to PPS U44 Details: no problem Status: Done. Konstantin


CERN_PPS (C5 Report on 31-Jan-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* gLite3.1.0 PPS Update 15 and 16 applied to 
   - AFS UI
   - CEs and WNs

* re-installation of DPM


  • 2007-12-30 : Task:AFS PPS UI update 16 Details:Status:Done. Farida

  • 2007-12-30 : Task:AFS PPS UI update 15 Details:Status:Done. Farida

  • 2007-12-29: Task:PPS-update 16 Details:All glite-WNs updated.Status:Done. Farida

  • 2007-12-23: Task:Reinstall DPM lxb2058 on SLC4. Details:Publishing DPM restarted.Status:Done. Farida

  • 2007-12-23: Task:PPS-update 14 . Details:CEs, Torque-serverupdated. Status:Done. Farida


CERN_PPS (C5 Report on 24-Jan-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* AFS UI upgraded to gLite3.1.0 PPS Update 13

* collaboration to pre-deployment tests of gLite 
   3.0.2 Update 44. Service tested:
   - FTS
   - LFC


CERN_PPS (C5 Report on 17-Jan-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Pre-deployment testing of  gLite3.1.0 PPS Update 13
   - top level BDII: new issue found in the distribution:  
          harcoded link to the FCR (bug opened)
   - due to a configuration issue during the pre-deployment 
     tests, the BDII was malfunctioning for 1 day on Monday, 
     which affected the submission of SAM tests.

* Site upgraded to gLite 3.0 PPS Update 43
   (FTS, MON+ e2emonit, WMS, LB)

* Support to CMS for test of Job Priority (in progress)

  • 2008-01-15: Task: gLite 3.1.0 PPS update 13, pps top BDII Details: The hard-coded web_url is fixed. A new bug #32655, FCR url hard-code. Status: Done. Lanxin

  • 2008-01-13: Task: FTS, WMSLB, MON+e2emonit to PPS U43 Details: no problem Status:Done. Konstantin

CERN_PPS (C5 Report on 10-Jan-08)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to report



CERN_PPS (C5 Report on 19-Dec-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to report



CERN_PPS (C5 Report on 13-Dec-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update11: In Progress
   - CEs : done
   - DPM: done
   - WNs: done
   - top BDII: done
   - UI: In progress
   - VOBOX: In Progress

  • 2007-12-13: Task:pps top BDII updated 11 . Details: updated. Status:Done. Lanxin

  • 2007-12-11: Task:PPS-AFS-UI updated 11 . Details:Tested and announced. Status:Done. Farida

  • 2007-12-11: Task:PPS-update 11 to site. Details: CEs,DPMs,glite-WNs updated.Status:Done. Farida


CERN_PPS (C5 Report on 06-Dec-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update10: In Progress
   - AFS UI: done
   - DPM: done
   - LFC: In progress

  • 2007-11-30: Task:Pre-deployment testing Details:CERN_PPS volunteered to cover oracle-based versions of LFC and DPM .Status:Done. Antonio

  • 2007-11-30: Task:Updated PPS-AFS-UI to update 10 . Details:Tested. updated.Status:Done. Farida

  • 2007-11-29: Task:Installed glite 3.1 DPM(MYSQL) PPS-update10 . Details: Tested. Find some errors during configuration. Reported to the release people. Applied manual fix.Status:Done. Farida

  • 2007-11-27: Task:Found problem in reinstallation of PPS machine SL4. Details: Error: GRUB Loading Stage2.... Status:Progress. Farida


 
CERN_PPS (C5 Report on 29-Nov-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Nothing to Report


CERN_PPS (C5 Report on 22-Nov-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.1 PPS-update09: Done
   - CEs, glite-WNs, site-BDII
* PPS AFS UI 3.1updated to 09
* Version v1.18-1 of CAs tested and deployed
* Set-up of lsf based CE lxb2090 implementing the new (uncertified) 
   version of the jobpriority configuration for Atlas and CMS: Done

  • 2007-11-22: Task:PPS AFS-UI updated to update09. Details:tested. updated.Status:Done. Farida

  • 2007-11-21: Task:Applied PPS-update09 . Details:All CEs, glite-WNs, site-BDII updated.Status:Done. Farida

  • 2007-11-14: Task:Applied fixed on AFS-UI for PPS for the VOMS cert. Details: Status:Done. Farida

  • 2007-11-14: Task:Setup LSF CE:lxb2090 for Job priority test for ATLAS and CMS. Details: glite 3.1 version installed. Status:Done. Farida

  • 2007-11-19: Task: Configured lcg-CE: lxb1937 for the Job priority test for CMS. Details:lxb1937 now support two VOs: ATLAS, CMS. Status: Done. Farida

CERN_PPS (C5 Report on 14-Nov-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.0 PPS-update42: Done
   - CEs, DPM, FTS, FTS, MONBOX+e2emonit, WMSLB

* High load on roc-wms, used as fail-over node for SAM Admins Page
   and Cern ROC certification

* Set-up of lsf based CE lxb2090 implementing the new (uncertified) 
   version of the jobpriority configuration for Atlas: In progress


  • 2007-11-14: Task:Setup LSF CE:lxb2090 for Job priority test for ATLAS. Details: glite 3.1 version installed. Status:Progress. Farida

  • 2007-11-14: Task:PPS-update 42. Details: CE, DPMs updated. Status:Done. Farida

  • 2007-11-14: Task: FTS, WMSLB, MON+e2emonit to PPS U42 Details: no problem Status:Done. Konstantin

  • 2007-11-13: Task: roc-wms update to PPS-U42 Details: no problem Status:Done. Lanxin

  • 2007-11-12: Task: roc-wms (#28929 in GGUS): system load is too high. Details: Fixed. Status:Done. Lanxin

  • 2007-11-08: Task: FTM: Problem report to GOC DB. Details: New node type 'FTM' requested (GGUS ticket #28795). Status:Done. Antonio

  • 2007-11-08: Task: MON, e2emon--> PPS-U41 Details: None. Status: Done. Konstantin

  • 2007-11-08: Task: FTS --> PPS-U41 Details: None. Status: Done. Konstantin

  • 2007-11-08: Task: FTM --> PPS-U41 Details: see next entry for the bug Status: Done. Konstantin

  • 2007-11-08: Task: FTM, FTS, WMSLB, MON pointed to CNAF PPS repo. Details: services pointed to CNAF PPS repo. Status: Done. Konstantin

  • 2007-11-08: Task: FTM reconfiguration problem. Details: see bug report #31100 Status: Progress. Konstantin


CERN_PPS (C5 Report on 08-Nov-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to glite 3.0 PPS-update41
   - Torque server done
   - AFS UI in progress
   - An issue in the distribution of the new certificate of the voms server
     was found during the upgrade of the WMSLB: the problem was reported
     to the ITR team and it triggered a new high priority patch to be prepared 
     for production and PPS

* Re-installation from scratch of old lsf-based CE
   in progress

* A new test Torque CE (lxb1937) was set up implementing the new (uncertified) 
   version of the jobpriority configuration for Atlas.
   The installation of the corresponding lsf test machine is in progress 


  • 2007-11-08: Task:Reinstallation of LSF-CE lxb2090. Details: The machine has old configuration and it has to be scratch. Status: Progress. Farida

  • 2007-11-07: Task: Job priority test on lxb1937(3.1 SLC4) Details:Installation done by scratch and tested. Status: Done. Farida

  • 2007-11-06: Task:Torque-server updated to PPS-update 41. Details: Update was fine. Status: Done. Farida

  • 2007-11-06: Task: PPS-AFS-UI 3.0 updated to 41. Details: Its ready but not yet announced. Still need some time to test. Status: Done. Farida

  • 2007-10-31: Task: LSF upgraded to 7.0 on lxb2087 Details:Latest version installed and configured. Status: Progress. Farida

  • 2007-11-01: Task: lcg-CE:lxb2006 SLC4 installed and tested. Details: Maui package was missing. Talked to the release group. Issue was resolved. Status: Done. Farida


CERN_PPS (C5 Report on 01-Nov-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Site upgraded to glite 3.1 PPS-update08 
   - AFS UI, WNs, site-BDII done
   - CE in progress

* roc-wms.cern.ch, used by SAM Admin's Page, updated to gLite 3.0.2 PPS-update41

* Security alerts triggered by e2emonit service: The root cause was an expired certificate 
   which was renewed on the node but not properly re-mapped in service internal directories
   Two bugs were opened in Savannah
    - https://savannah.cern.ch/bugs/index.php?30735 against e2emonit
    - https://savannah.cern.ch/bugs/index.php?30736 against netmon2rgma
   


  • 2007-11-01: Task: PPS-AFS-UI 3.1 updated to 08 Details: Tested and BROADCAST was send. Status: Done. Farida

  • 2007-10-31: Task: PPS-AFS-UI updated to update08 Details: Status: Progress. Farida

  • 2007-10-31: Task:PPS-update08 Details: The CE will be scratched.DPM,glite-WNs updated. Status: Progress. Farida

  • 2007-10-30: Task:Debug LSF-CE. Details: Gip timed out. under investigation. Status: Progress. Farida

  • 2007-10-30(1): Task: roc-wms Details: roc-wmc.cern.ch is updated to PPS_update 41. Status: Done. Lanxin


CERN_PPS (C5 Report on 25-Oct-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Site upgraded to glite 3.1 PPS-update 07 (UI, WNs, site-BDII)
  - Site BDII: we observed an issue not strictly related to YAIM configuration but dealing with information "hardcoded" in the site BDII (the normal hack you have to do in order to publish castor srms. As the intelligence of setting mdo-vo-name- was moved down from the site bdii to the gip itself, the ldif files we used before to hardcode info about  the srms were not correct anymore we had to change our files as shown in the diff below: This information should be passed to CERN-PROD
  - the site BDII was stopped and not released to production yet in order for a configuration bug to be fixed
* Old versions of the AFS gLite User Interface removed. No complaints received from users 

  • 2007-10-18(2): Task: AFS UI Details: Cleanup of old AFS directory done and broadcast to users sent. Status: Done. Antonio

  • 2007-10-18(1): Task: Firewall/access administration Details: GD_FTM profile applied to the service node. Konstantin, Steve and myself access the node by default. Status: Done. Antonio


CERN_PPS (Section and C5 Report on 18-Oct-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Site upgraded to glite 3.0 PPS-update 40 (CE, DPM, WNs)
* top-level BDII upgraded to glite 3.1 PPS-update07
* roc-wms.cern.ch, WMS for CERN ROC and SAMAP reinstalled with latest PPS update available 
* Two bugs (one critical) opened against BDII in gLite3.1.0 Update07

  • 2007-10-04: Task:Applied glite 3.0 PPS-update 40 Details:All CEs, DPMs,glite-WNs updated. Status: Done. Farida

  • 2007-10-16: Task: Reinstall PPS site-BDII to glite 3.1 SL4 Details:Applied PPS-update 07 to lxb2002 .For short time the service was down. Status: Done. Farida

  • 2007-10-17(3): Task: roc-wms Details: re-installed roc-wms.cern.ch with the latest version of pps-WMSLB Status: Done. Lanxin

  • 2007-10-17(2): Task: gLite BDII 3.1 with SLC4(patch#1385) Details: open a bug #30486 about web_url hard-core to lcg bdii production in glite-info-update-endpoints Status: Done. Lanxin

  • 2007-10-17(1): Task: pps-bdii Details: Alias switched from lxn1187 to lxb2055. GOC db updated. Broadcast sent. Status: Done. Antonio


CERN_PPS (Section and C5 Report on 11-Oct-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* New service File Transfer Manager was installed and configured lxb2039.cern.ch. 
   Feedback provided to integration and release team.
* PPS AFS UI 3.1updated to 06
* Version v1.17-1 of CAs tested and deployed
* SAM CE test submission from the UI(lxb2086) to 
   roc-wms failed due to a bug in the UI. Reported
* New version of the kernel deployed on all SLC3 nodes

  • 2007-10-10: Task: New service FTM. Details: New service File Transfer Manager was installed and configured lxb2039.cern.ch. Feedback provided to integration and release team. Status: Done. Konstantin

  • 2007-10-10: Task: PPS AFS UI 3.1updated to 06 (Test) Details:Spend some time on update of pps afs UI 3.1 and made it work. Status: Done. Farida

  • 2007-10-10: Task: Kernel upgraded. Details: Kernel upgraded on site-BDII, CEs, DPMs, glite-WNs ,Torque-Server to 2.4.21-52. Status: Done. Farida

  • 2007-10-10: Task: Kernel upgrade on FTS/WMS/MON_e2emonit. Details: Kernel upgraded to 2.4.21-52 on the hosts: pps-wms, pps-fts, pps-mon. Status: Done. Konstantin

  • 2007-10-09: Task: CA (v1.17-1) updated and tested. Details: CEs, DPMs,glite-WNs updated. Status: Done. Farida

  • 2007-10-09(1): Task: Found a bug(#30200) of yaim-4.0.0-13 when configuring gLite-UI 3.1 with SLC4 Details: ROC SAM CE test submission from the UI(lxb2086) to roc-wms faild due to the bug Status: Fixed. Lanxin

  • 2007-10-02(1): Task: PPS-glite-BDII 3.1 Details: Installed PPS-glite-BDII on lxb2055 with SLC4 (GGUS #27311), glite-version rpm is missing in the repository, used new BDII node type to configure it. So the new BDII works . Status: Done. Lanxin


CERN_PPS (Section and C5 Report on 27-Sep-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* gLite 3.0.2 PPS update 40 applied to top-level BDII(pps-bdii.cern.ch) 
   within PPS pre-deployment testing
* SAM tests results not available for lxb2002. Node correctly registered 
   in SAM DB. GGUS ticket #27041 opened against APROC (bdii sensor)

  • 2007-09-26(2): Task: Web site Details: CERN_PPS web Page www.cern.ch/pps/index.php?dir=./PPSsites/CERN_PPS/ updated with links to usful tools (CERN_PPS page on GOC DB, LCG Node status page, worklog) Status: Done. Antonio

  • 2007-09-26(1): Task: PPS update 40 Details: in CERN pps top BDII(pps-bdii.cern.ch) Status: Done. Lanxin

  • 2007-09-24(1): Task: AFS UI Details: information about CERN AFS UI should be published in CESGA wiki (private, cern area). The information is now linked from the web site and the public wiki. Status: Done. Antonio

  • 2007-09-21(2): Task: AFS UI Details: group aretico:uiadmin created including diana, fnaz, jnovak, aretico, dimou, maart . This group is given writing access to the afs Ui directory. Status: Done. Antonio

  • 2007-09-21(1): Task: site bdii Details: SAM tests not available for lxb2002. Node correctly registered in SAM DB. GGUS ticket #27041 opened against APROC (bdii sensor) Status: Done. Antonio


CERN_PPS (Section and C5 Report on 20-Sep-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Site upgraded to gLite 3.1.0 PPS-Update 07
* WMS was replaced with new one
* CERN_PPS web Page www.cern.ch/pps/index.php?dir=./PPSsites/CERN_PPS/ updated 
with links to various internal monitoring systems (Nagios, Ganglia, Yvan's tool)
* wrong script name to source for UI configuration fixed in UI documentation on wiki
* Two bugs raised against integration and voms within resolution of ticket #26968

  • 2007-09-18: Task: Documentation Details: Two bugs raised against integration and voms within resolution of ticket #26968 Status: Done. Antonio

  • 2007-09-17: Task: New WMS installed. Details: New WMS was installed from scratch (lxb2179) due to lack of space on partition for SandBoxes on the previous one (lxb2039). Status: Done . Konstantin

  • 2007-09-14: Task: Documentation Details: wrong script name to source for UI configuration fixed in UI documentation on wiki Status: Done. Antonio

  • 2007-09-14: Task: Web site Details: CERN_PPS web Page www.cern.ch/pps/index.php?dir=./PPSsites/CERN_PPS/ updated with links to various internal monitoring systems (Nagios, Ganglia, Yvan's tool) Status: Done. Antonio


CERN_PPS (Section and C5 Report on 13-Sep-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Site upgraded to gLite 3.0.2 PPS-Update 39
* patch #1285 for gLite3.0 dealing with proxy-renewal on WMS, 
   to be released in production with gLite 3.0 Update34 
   was applied and tested on CERN_PPS WMS out of the standard release process. 
   Green light given for release.
* Maintenance interventions on WNs and troubleshooting on CEs 


  • 2007-09-13: Task: To fast-track a patch for WMS to PROD.Details: We were asked by SA3 to fast-track a patch to PROD by "quickly" testing it on CERN_PPS WMS. The patch was related to proxy-renewal on WMS. Patch was applied and tested. Green light given for release. Status: Done. Konstantin

  • 2007-09-10: Task:Reinstalled glite-WN lxb1934. Details: old installation.: Status:Done .Farida

  • 2007-09-07: Task: CEs, DPMs, glite-WNs upgraded to PPS-update 39.Details: Status:Done. .Farida

  • 2007-09-05: Task:Troubleshooting CEs, failing SAM JS test. Details: Debug of lxb1926 (Torque server) BLParser restarted. Status: Done . Farida

  • 2007-08-27: Task: WMS from U38 - inst/conf/tested. Details: New WMS instance (lxb2039.cern.ch) installed/configured/tested from scratch. Required by PPS U38. Problems found with GridFTP - bug reports submitted/updated (#18497, #29031) Status: Done . Konstantin


CERN_PPS (Section and C5 Report on 23-Aug-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* Upgrade of the site to gLite 3.0.2 PPS-Update 38 in progress
* Reconfiguration of the FTS web service in order to use Oracle 
  DB writer account
* Support for SAMAP. Upon request of SAMAP, the SAM client used 
   for CERN ROC now submits also to all uncertified PPS sites 


  • 2007-08-23: Task: WMS from U38. Details: Started with installation/conf. of new WMS instance (lxb2039.cern.ch). Required by PPS U38. Status: Done . Konstantin

  • 2007-08-22: Task: FTS web-service was reconfigured to use Oracle DB writer account. Details: FTS web service was reconfigured to use Oracle DB writer account. Status: Done . Konstantin

  • 2007-08-20: Task:Setup new SAM submission for PPS uncertified sites of all regions on lxb2086 Details: This submission runs each 2 hours for dteam vo Status: Done . Lanxin

  • 2007-08-16: Task:Setup new lcg-CE. Details: network scan, preliminary to firewall opening, requested for lxb1914 Status: In Progress . Antonio


CERN_PPS (Section and C5 Report on 16-Aug-07)

CERN GRID Pre-Production Site (CERN_PPS):
-----------------------------------------
* The Involvement of CERN_PPS in the official CA deployment procedure 
  is now documented. As we were alreday "silently" doing CERN_PPS will 
  test the installation of the CA distribution from EUGRIDPMA in parallel 
  with the development of up-to-date CA tests in SAM
  details about the procedure in
  https://wiki.egi.eu/wiki/EGI_IGTF_Release_Process
  details about CERN_PPS involvment in
  https://twiki.cern.ch/twiki/bin/view/LCG/PPS_CA_Test
  A node will be reserved for testing the installation from scratch
* Decommissioning of gLite CE completed
* CA update; All services and SAM client updated
* New lcg-CE installed. It is meant to replace an old server. 

 

  • 2007-08-16: Task: Fixed GGUS Ticket#25516 for lcg-CE Details: SAM Job submission test was not passed. Status: Done . Farida

  • 2007-08-16: Task: Reconfigured site-BDII to published BNL srmv2 point correctly. Details: Status: Done . Farida

  • 2007-08-16: Task: glite-CEs removed from infomation system. Details: glite-CEs are removed from site-BDII. Status: Done . Farida

  • 2007-08-16: Task:Setup new lcg-CE lxb1914. Details: Installed & configured lcg-CE lxb1914 for CERN-PPS and tested. Status: Done . Farida

  • 2007-08-16: Task:latest CA Test. Details: CA updated to 1.16 Status: Done . Farida

  • 2007-08-15(1): Task: Testing. Details: SAM sensor upgraded to 1.3.0-2 on CERN ROC SAM client Status: Done . Lanxin

  • 2007-08-10(1): Task: Testing. Details: SAM sensor was updated to 1.2-4 on lxb2086(CERN ROC SAM client) due to new version CA-1.16-1 Status: Done . Lanxin

  • 2007-08-10(1): Task: Testing. Details: lcg-CA was updated to 1.16-1 on lxb2086 (gLite-UI3.1 and SAM client of CERN ROC) Status: Done . Lanxin

  • 2007-08-13(1): Task: Testing. Details: Involvement of CERN_PPS in CA deployment procedure documented www.cern.ch/pps/index.php?dir=./panel/SVC/& Status: Done . Antonio


CERN_PPS (Section and C5 Report on 09-Aug-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
* hardware failure on pps-wms
  The service was replaced with a new node
  The UI was reconfigured to use a different default (CNAF)
* PPS CERN AFS UI (gLite 3.0): gLite 3.0 UPDATE 37 
   now available
* decommissioning of gLiteCEs started
 

  • 2007-08-07(1): Task: hardware failure on lxb2070 (pps-wms).Details: Service to be reinstalled on lxb2176 re-installed with SLC3, access fixed, the dns alias pps-wms now points to this node Status: In Progress . Antonio

  • 2007-08-06(2): Task: hardware failure on lxb2070 (pps-wms). Details: hardware failure on lxb2070 (pps-wms). Service to be reinstalled on lxb2176 Status: In Progress . Antonio

  • 2007-08-06(1): Task: Documentation. Details: links to afs UI (CERN) changed in the wiki Status: Done . Antonio


CERN_PPS (Section and C5 Report on 02-Aug-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
* Site updated to PPS-update 37
* Torque-server upgraded to 2.1.8.
*  Found bug#28456 in cronjob for DPM (that creates "generated" directories for all VOs on DPM).
* PPS Site-BDII is configured to publish SARA-SRM2 .
 

  • 2007-07-19(1): Task: Submitted bug#28456 found on DPM. Details: cronjob failed to create default "generated" directories for each VO on DPM with octal number greater then 07. Applied temporary fix on lxb2058,lxb2088 for PPS. Status: Done . Farida

  • 2007-07-19(1): Task: PPS Torque-server upgraded to 2.1.8 Details:Since old version had many problems. Job stucked in queue forever. Status: Done . Farida

  • 2007-07-19(1): Task: Site updated to PPS-update 37 Details: CEs,WNs,DPMs updated. Status: Done . Farida


  • 2007-07-26: Task: PPS U37 on FTS and MONBox_e2emonitDetails: No problems. Status: Done . Konstantin

  • 2007-07-25: Task: Helped to debug MyProxy server related problem on FTS instance at CNAFDetails: Now it seems to be working for CNAF. Status: Done . Konstantin

  • 2007-07-25: Task: Port changes for CE in site-BDII. Requested by FIODetails:None Status: Done . Farida

  • 2007-07-25: Task: Made request to monitor BDII for PPS by SAM.Details: Currently SAM is not monitoring the site-BDII and top-level BDII for all the PPS. Only turn ON the monitoring in GOC DB is not enough. Status: Progress . Farida

  • 2007-07-25: Task: PPS-update 36 Details: CEs,DPMs,WNs updated. Status: Done. Farida


CERN_PPS (Section and C5 Report on 19-Jul-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
* Support for SAMAP / Cern ROC
   - new SAM submission chain (WMS-BDII-UI-SAM Client) up and running
   - regular submission on behalf of 'dteam' to PPS/PROD Certified/Uncertified sites in CERN ROC 
   - documentation available in https://twiki.cern.ch/twiki/bin/view/LCG/CERNROC
* New  Diligent voms server certificate deployed on the site

 

  • 2007-07-17: Task: Support for SAMAP and CERN ROC, a new SAM submission chain Details: Finish installation and configuration of SAM client for CERN ROC(lxb2086) which test certified and uncertified sites of production and PPS. The new SAM submission chain is created. And wrote twiki page Status: Done. Lanxin

  • 2007-07-17: Task: Support for SAMAP and CERN ROC, a new SAM submission chain Details: Finish deploying glite-UI 3.1 with SLC4(lxb2086) Status: Done. Lanxin

  • 2007-07-17: Task: Support for SAMAP and CERN ROC, a new SAM submission chain Details: Finish deploying ROC glite-WMS server(lxb2059) which support dteam and ops VOs Status: Done. Lanxin

  • 2007-07-11: Task: Diligent VO: Will do some tests for data challenge at CERN-PPS. Details: Diligent new voms server certificate placed on test cluster. First job submission test was successful. Status: Done. Farida,Antonio,Konstantin


CERN_PPS (Section and C5 Report on 12-Jul-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
* Upgrade of the site to gLite 3.0.2 PPS-UPDATE 35 
   (New YAIM) in progress
    - CEs, DPMs, LFC, lcg-RB, BDII, FTS done
    - yaim bugs: #25667, #27984.
* AFS UI released (Update 02)
   - first user immediately complained for change of location of 
     setenv script (although advertised in the broadcast)
   - need for the CERN-PROD (ROC) to advertise the change more noisily
     before going to production. Also a longer transition period, with symlinks 
     available in parallel would help 
* Job submission was failing for sgm users. We needed to add more pool accounts for 
   sgm users on CE and Torque-server.
   - This solution is correct from the site point of view.
   - The reason why the problem happened, though, is because the 
      'lcgadmin' role in dteam is unreasonably crowded (GGUS#22814)
   - The role needs to be cleaned up (action for OCC and ROCs)
 

  • 2007-07-12: Task: glite 3.0.2 PPS Update 35 Details: FTS - no problems. Status: Done. Konstantin

  • 2007-07-12: Task: glite 3.0.2 PPS Update 35 Details: Upgraded on pps VOBOX (lxn1186.cern.ch), lcg-RB (lxb2085.cern.ch), glite-LFC_mysql (lxb2038.cern.ch) Reconfigure and restart. Status: Done. Rolandas

  • 2007-07-11: Task: Support for SAMAP and CERN ROC, a new SAM submission chain Details: deploy and configure ROC WMS+LB(lxb2059) Status: in progress. Lanxin

  • 2007-07-10: Task: glite-yaim upgraded to 3.1.1-8 Details: Upgraded on pps top BDII(pps-bdii.cern.ch) Reconfigure and restart top BDII service. Status: Done. Lanxin

  • 2007-07-10: Task: glite-yaim upgraded to 3.1.1-8 Details: Upgraded on glite-CE, DPM. Reconfigure the nodes. Status: Done. Farida

  • 2007-07-10: Task: Troubleshooting on lcg-CE. Details: Job submission was failing for sgm users. Add more pool accounts for sgm users on CE and Torque-server. Tested by Konstantin with lcg-admin role, successfull. Status: Done. Farida

  • 2007-07-10: Task: CA upgraded to 1.15 Details: Upgraded CEs, DPMs, glite-WNs to the latest version. Did some basic tests on DPM which was successfull. Status: Done. Farida

  • 2007-07-10: Task: AFS UI Details: Wrong DN for lcg-voms set in vomses file. Manually fixed Status: Done. Antonio

  • 2007-07-07: Task: Site operations/ Details: Host certificates requested and installed on 8 nodes Status: Done. Antonio

  • 2007-07-06: Task: Support for SAMAP and CERN ROC, a new SAM submission chain Details: Finish deploying ROC top BDII(lxb2069) which contains certified and uncertified sites Status: Done. Lanxin


CERN_PPS (C5 Report on 05-Jul-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * Upgrade of the site to gLite 3.0.2 PPS-UPDATE 34 
      (Security patche on DPM) done
        - 3 DPMs upgraded and successful basic testing 
   * Support for SAM Admin's Page (installation of "special" BDII + WMS)
      in progress.
   * CMS testing SLC4 worker nodes through CERN_PPS . CMS is mostly 
      interested in the cluster maintained by FIO. Currently waiting for FIO to 
      upgrade the CEs to the last version of middleware.
   * CERN_PPS "officially" stopped the submission of SAM tests to PPS
      The instance will now be used as a pilot to tune FTS testing in PPS



CERN_PPS (Section Report on 05-Jul-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * Upgrade of the site to gLite 3.0.2 PPS-UPDATE 34 
      (Security patche on DPM) done
        - 3 DPMs upgraded and basic testing successful
   * Kernel upgraded on 19 nodes
   * We gave access yo one of our WNs to Ulrich, from FIO, in order to help 
      with the upgrade of the CERN-PROD farm to Natively compiled SLC4 WNs
   * Support for SAM Admin's Page (installation of "special" BDII + WMS)
      in progress.
   * CMS testing SLC4 worker nodes through CERN_PPS . CMS is mostly 
      interested in the cluster maintained by FIO. Currently waiting for FIO to 
      upgrade the CEs to the last version of middleware.
   * CERN_PPS "officially" stopped the submission of SAM tests to PPS
      The instance will now be used as a pilot to tune FTS testing in PPS
 

  • 2007-06-21: Task: Troubleshooting glite-CE(PPS). Details: Failing SAM job submission test. Status: Progress. Farida

  • 2007-07-04: Task: Debug lcg-CE. Details: cms user got problem while mapping to the pool account. Created more pool accounts for the sgm user. Status: Done. Farida

  • 2007-07-04: Task: Upgraded kernel to kernel#2.4.21-50.EL.cern at CERN-PPS. Details: manualy upgraded kernel on almost 19 machines. Status: Done. Farida

  • 2007-07-04: Task: Help provided to Ulrich (FIO) in upgrade glite-WN SLC4 3.1. Details: Gave interactive access to one of machine. Status: Progress. Farida

  • 2007-07-02: Task:Upgraded DPMs without any problem. Details: PPS-UPDATE34 applied to lxb2058,lxb2088,lxb1914. I did some initial tests which seemed fine. Status: Done. Farida

  • 2007-06-21: Task: Support for SAMAP and CERN ROC Details: Lxb2059/69/86 reinstalled, Status: Done. Antonio


CERN_PPS (C5 Report on 28-Jun-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * Upgrade of the site to gLite 3.0.2 PPS-UPDATE 33 in progress
       * CE, DPM, BDII done
       * no major issues except short downtime due to GLUE schema 
          upgrade
   * CERN_PPS volunteered to maintain two services (BDII + WMS) 
      specially configured to work as failover nodes for SAMAP (SAM Admin's Page)
      service. Lanxin will be te service administrator. The chain will be completed
      by a brand new SAM client and UI meant to submit jobs to sites in
      Cern ROC scope. In progress
   * The new Nagios-based monitoring tool is now monitoring PPS sites:
      James Casey, and Farida on behalf of the Monitoring Working Group, are looking at 
      results of CERN_PPS in order to fine tune the tool, in comparison with alarms
      raised from CERN 
   * SRMv2 testing: Flavia found out a general issue in the information published by SRMs
      DPMs at CERN_PPS were re-configured accordingly
   


  • 2007-06-26: Task: PPS-BDII upgraded to pps-update 33. Details:No problems found Status: Done. Lanxin

  • 2007-06-21: Task: CEs , DPM, site-BDII upgraded to pps-update 33. Details:No major problems found except due to schema upgrade the services got down for a short time. Status: Done. Farida

  • 2007-06-21: Task: Support fpr SAMAP and CERN ROC Details: A new SAM submission chain (SAM client, UI, WMS, BDII) will be maintained by lanxin in CERN_PPS. WMS and BDII will contain unceritifed sites and will work as fail-over for the SAMAP client, regular SAM submission to uncertified sites will be done by the CERN ROC Status: In progress. Antonio, Lanxin

  • 2007-06-21: Task: "Nagios" a new monitoring tool started to monitor the CERN-PPS site. Details: none Status: Done. Farida

  • 2007-06-21: Task: Start using Alarms in result of SAM failure & compare the failure raised by the montoring tool "Nagios". Details: Activated Site Alert Notification for all the services at CERN-PPS and compare the failure results with "Nagios" a new monitoring tool. Status: Done. Farida

  • 2007-06-22: Task: Yaim bug: Incorrect service type was publised by the DPM. Details:Applied manual configuration changes to all the DPMs at CERN-PPS in order to publish correct service type. Status: Done. Farida

  • 2007-06-22: Task: Filled RC pre-report for CERN-PPS site on CIC portal . Details:none Status: Done. Farida

  • 2007-06-22(2): Task: SRM v2 testing. Details:srm service type needs to be changed in lxb2058 and lxb2088. I analysed the YAIM function config_gip to find the place where to do the change. Status: In progress. Antonio

  • 2007-06-22(1): Task: Reports. Details:CERN_PPS pre-report (CIC portal) filled. Status: Done. Antonio


CERN_PPS (C5 Report on 21-Jun-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * Upgrade of the site to gLite 3.0.2 PPS-UPDATE 32 completed
   * Upgrade to gLite 3.1.0 PPS-U01 (UI)
      During the installation we noticed that 

      1) The directory $INSTALL_ROOT/etc was removed and its 
          content put in $INSTALL_ROOT/external/etc . 
          As a major consequence for the user
          the location of the environment script has changed
      2) The name of the environment script was changed from grid_env.(c)sh 
          to grid-env.(c)sh, 
      
      Both these changes are likely, in our opinion, to create surprise and discomfort in 
      the user community.
Therefore we opened two critical bugs to track these issues. - https://savannah.cern.ch/bugs/index.php?27361 - https://savannah.cern.ch/bugs/index.php?27362

  • 2007-06-17: Task: Job submission failed on lcg-CE . Details:After debug the problem was due the wrong time synchronization between CE and WN. Status: Done. Farida

  • 2007-06-20: Task: gLite3.1.0 pps-U01 . Details: UI

  • 2007-06-17: Task: pps-update 32 . Details: CEs, SEs, site-BD updated. Status: Done. Farida

  • 2007-06-17: Task: U32 on WMSLB, FTS, MON_e2emonit. Details: no problems during update. Status: Done. Konstantin


CERN_PPS (C5 Report on 14-Jun-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * Upgrade of the site to gLite 3.0.2 PPS-UPDATE 32 in progress
   * LHCb user requested and obtained interactive account on one of the SLC4 WNs
   * Temporary un-availability of FTS due to Oracle password expiration on Monday


  • 2007-06-14: Task: afs administration Details: I created a afs group aretico:cern-pps including all administrators in CERN_PPS. This group was firstly applied in the yaim-server directory Status:Done. Antonio

  • 2007-06-11: Task: Installed DPM (MYSQL) Details:This will be a test machine where certification people will test some new functionality of DPM. Access will be given. Status:Done. Farida

  • 2007-06-11: Task: Updates of Preproduction report failed. Details: I found some problems on CIC Portal in updating of Preproduction report for CERN_PPS site. Reported the problem. Status:Progress. Farida

  • 2007-06-11: Task: Password for Oracle DB account 'lcg_egee_pps_fts' has expired. FTS was not working. Details: Oracle team was notified - ticket CT433723. Then figured out that the password can be reset by user. So, the password was reset an now the service is back to work. However, there are two accounts used by FTS 'lcg_egee_pps_fts' and 'lcg_egee_pps_fts_w' - so, the password for "writer" account did not expire and remained the same (site-info.def changed accordingly, service reconfigured (2 accounts, 2 passwords)). The password for "writer" account may expire in the near future. Status: Done. Konstantin

CERN_PPS (C5 Report on 7-Jun-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * UPDATE 31 successfully applied to all services
   * Helping ATLAS with SRM v2 tests via CERN_PPS FTS2.0:
      - First successful transfer with FTS 2.0 for ATLAS done through FTS v2 at CERN_PPS
      - Changes were needed in the information published by site BDII
   * Helping experiments with SLC4 WN testing
   * After SRMv2 service was introduced in SAM, we cannot see anymore the new strorage 
      elements in the display (they were previously reported as SEs and SRMs). We are investigating 
      on the issue in order to submit an appropriate ticket

  • 2007-06-06: Task: PPS-BDII. Details: PPS-BDII was updated to PPS-update 31 Status: Done. Lanxin

  • 2007-06-04: Task: Fixed site-BDII. Details: pps site-BDII was fixed to publish the correct information for SRM v2: BNL-V22 Status: Done. Farida

  • 2007-06-31: Task: CEs , site-BDII, DPMs updated. Details: PPS-update-31 applied without any problems. Status: Done. Farida

  • 2007-06-01: Task: Reinstall glite-CE. Details: Reason: Old installation. Status: Done. Farida

  • 2007-06-04: Task: U31on WMSLB, FTS, MON_e2emonit. Details: no problems during update. Status: Done. Konstantin

CERN_PPS (C5 Report on 31-May-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * SRM v2 Testing in progress
   * Testing native SLC4 WN 3.1 by the experiments in progress.  
    Interactive access to one of the SLC4 3.1 WNs was granted to the ATLAS and LHCb for performing tests.
   * Helping ATLAS with SRM v2 tests via CERN_PPS FTS2.0.

* 2007-06-31: Task: VO tests for SLC4 WN 3.1 Details: Gave interactive access to the users of ATLAS,LHCb Status: Done. Farida

  • 2007-05-30: Task: Helping ATLAS with SRM v2 tests via CERN_PPS FTS2.0 Details: required additional channels were set up in FTS. Problems are mostly with SRM v2 endpoints configurations and info published in IS. Status: Done. Konstantin

  • 2007-05-24: Task: SRM v2 tests Details: site bdii re-configured to fix the information published for LBNL-BESTMAN (see GGUS #22339) Status: Done. Antonio


CERN_PPS (C5 Report on 24-May-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * SRM v2 Testing in progress: 
     - Active testing of channel CERN-PROD --> BNL. 
       Serious configuration problems found on dCache at BNL 
       (This problem is common to all dCache SRMv2 installations and 
         needs advise by developer,currently on leaves).
 

CERN_PPS (Section Report on 24-May-07)

No section meeting


  • 2007-05-10: Task: Reporting Details: Section and C5 report written in worklog Status: Done. Antonio


CERN_PPS (C5 Report on 10-May-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
   * SRM v2 Testing in progress: 
     - PPS Site-BDII is now configured to publish srmv2.2 end points. 
     - lcg-CE reconfigured: srmv2.2 end points added as close SEs.
     - Requirements received form Atlas for FTS channel configuration
   * FTS2 migration in progress: 
     - DB schema migration problem due to wired insufficient privileges to create indexes. Took some time to 
       resolve the problem. Oracle support missed the submitted ticket. After running the DB migration script for
       the second time there is a problem that transfer jobs stay in "Submitted" state. Investigating the problem.
   * Upgrade to PPS UPDATE 29
     - AFS UI, CE and DPM done
     - VOBOX couldn't be upgraded because of dependency problem. bug reported


CERN_PPS (Section Report on 10-May-07)

   * SRM v2 Testing in progress: 
     - PPS Site-BDII is now configured to publish srmv2.2 end points. 
     - lcg-CE reconfigured: srmv2.2 end points added as close SEs.
     - Requirements received form Atlas for FTS channel configuration
   * FTS2 migration in progress: 
     - DB schema migration problem due to wired insufficient privileges to create indexes. Took some time to 
       resolve the problem. Oracle support missed the submitted ticket. After running the DB migration script for
       the second time there is a problem that transfer jobs stay in "Submitted" state. Investigating the problem.
   * Upgrade to PPS UPDATE 29
     - AFS UI, CE and DPM done



  • 2007-05-10: Task:CERN_PPS Site-BDII Details:PPS Site-BDII is now configured to publish srmv2.2 end points. Status: Done. Farida

  • 2007-05-10: Task: FTS2 migration. Details: DB schema migration problem due to wired insufficient privileges to create indexes. Took some time to resolve the problem. Oracle support missed the submitted ticket. After running the DB migration script for the second time there is a problem that transfer jobs stay in "Submitted" state. Investigating the problem. Status: Progress. Konstantin

  • 2007-05-10: Task:lcg-CE Details:Reconfigured and added srmv2.2 end points as close SEs. Status: Done. Farida

  • 2007-05-10: Task:glite-CE Details: The reconfiguration of glite-CE (after update 28), the condor did not started and test jobs failed to submit. After debugging seemed the communication problem between the WMS and CE. the problem was resolved after some time.. Status: Done. Farida

  • 2007-05-10: Task: CEs , DPM Details: PPS-update 29. Status: Done. Farida

  • 2007-05-08: Task: AFS UI Details: Installed UPDATE 29. Status: Done. Diana

  • 2007-05-04: Task: Host Certificate Details: replaced for lxb1764 lxb1923. Status: Done. Antonio

  • 2007-05-03: Task: Site BDII Details: bdii.update manually re-configured to fix a srm publishing issue, site-info.def on the machine needs to be configured accordingly. Status: In progress. Antonio

  • 2007-05-03: Task: Reporting Details: C5 report created, Section abd C5 report written in worklog Status: Done. Antonio

  • 2007-05-03: Task: Reporting Details: Section report created. Status: Done . Konstantin


CERN_PPS (C5 Report on 3-May-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================
* PPS Update 28: in progress
   - CEs DPM, FTS Done

* Re-configuration of site BDII to publish SRM end-points 
  for SRMv2 testing in progress

* Malfunctioning of LB server affecting job submission via WMS pps-wms.cern.ch 
  - Bug submitted https://savannah.cern.ch/bugs/?25976. 
  - Access to pps-wms.cern.ch given to a developer. 
  - After debugging: #25976 updated with quick workaround. 
  - "Fix is in CVS. Patch will follow shortly." 
  - Workaround apllied and job submission restored

* Admins in CERN_PPS registered to the new "site alert notification" tool in the CIC Portal.
   So far no notification were received (because the site never failed)

* Debugging activity on lcg-CE (submission to SLC4 WNs) and DPM (falures in Gridftp)

* E2EMONIT (R-GMA nwtwork monitoring - High ping rate from UKI-LT2-IC-HEP-PPS. Details: 
   CERN firewall raised alerts. UKI-LT2-IC-HEP-PPS site admins and e2emonit developers contacted. 
   Issue solved.


CERN_PPS (Section Report on 3-May-07)
* FTS2.0 - installation bug after PPS U28. Details: Bug submitted 
  https://savannah.cern.ch/bugs/index.php?26048.

* CEs, DPM updated to the PPS - update 28. Status:

* Firewall profiles. Details: Two nodes for HeC registered, respectively, as 
   (GD_WMS+ GD_LFC+GD_BDII) and GD_VOMS. Security scan done. 
   External firewall opened.

* Job submission was not possible via WMS pps-wms.cern.ch - LB server was malfunctioning. 
  Details: Bug submitted https://savannah.cern.ch/bugs/?25976. 
  Access to pps-wms.cern.ch granted to a developer. -- After debugging: #25976 updated with quick 
  workaround. "Fix is in CVS. Patch will follow shortly." 
  
* Test CIC portal site alert notification in case of SAM failure. Status: Progress.

* DPM Host certificate was expired. New certificate was issued. Status: Done.

* Debugging of the lcg-CE(problems with the submission to one of slc4 batch
node) Status: Progress.

* Gridftp cmd failed on DPM (CERN-PROD). GGUS ticket is submitted. Status: Done.

* E2EMONIT - High ping rate from UKI-LT2-IC-HEP-PPS. Details: CERN firewall raised alerts. 
  UKI-LT2-IC-HEP-PPS site admins and e2emonit developers contacted. 
   Issue solved. Status: Done.

  • 2007-05-03: Task: CEs, DPM updated to the PPS- update 28. Status:Done . Farida

  • 2007-05-02: Task: Firewall profiles Details: Two nodes for HeC registered, respectively, as (GD_WMS+ GD_LFC+GD_BDII) and GD_VOMS. Konstantin has now access to all these service type. Waiting for security to scan the machine before authorizing external connectivity. Status: Progress . Antonio

  • 2007-05-02: Task: Job submission is not possible via WMS pps-wms.cern.ch - LB server is malfunctioning. Details: Bug submitted https://savannah.cern.ch/bugs/?25976. Access to pps-wms.cern.ch is given to a developer. -- After debugging: #25976 updated with quick workaround. "Fix is in CVS. Patch will follow shortly." Status: Done. Konstantin

  • 2007-05-02: Task: Test CIC portal site alert notification in case of SAM failure. Status: Progress . Farida

  • 2007-05-02: Task:DPM Host certificate was expired. New certificate was issued. Status: Done . Farida

  • 2007-04-27: Task: Debugging of the lcg-CE(problems with the submission to one of slc4 batch node) Status: Progress . Farida

  • 2007-04-27: Task: Gridftp cmd failed on DPM (CERN-PROD). GGUS ticket is submitted. Status: Done . Farida

  • 2007-04-26: Task: E2EMONIT - High ping rate from UKI-LT2-IC-HEP-PPS. Details: CERN firewall raised alerts. UKI-LT2-IC-HEP-PPS site admins and e2emonit developers contacted. Status: Done . Konstantin


CERN_PPS (C5 Report on 26-Apr-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================

   * Many failures were observed for SAM replica management. They were due to a bad configuration of the PPS SAM client happened after its upgrade. Now all the CEs are "green"

   * Atlas started looking into SLC4 WNs. Interactive access to one of the node was granted to SImone Campana and Rod Walker at this purpose

   * We are suffering a lot because of two bugs in the LB service, which force us to restart the service from time to time (this week it happened twice). A brand new LB diagnostic was installed in PPS with Update26, so we offered to WMS developer to access our sever to debug. No reply came yet. 

   * FTS server was fixed and put back to production. It is ready for testing FTS2.0 (currently in Certification)

  

CERN_PPS (Section Report on 26-Apr-07)

   * FTS server was fixed and put back to production.   It is ready for testing FTS2.0 (currently in Certification)

   * An issue was found in the KickStart based PXE installation used by GD. Workaround in place.

   * Many failures were observed for SAM replica management. They were due to a bad configuration of the PPS SAM client happened after its upgrade. Now all the CEs are "green"

   * Atlas started looking into SLC4 WNs. Interactive access to one of the node was granted to SImone Campana and Rod Walker at this purpose

   * Still waiting for FIO to give us a timeline for the delivery of ce110 and ce111. These CEs are meant to give access to a test cluster installed with SLC4 WNs. We warned FIO about difficulties in the upgrade. So FIO is currently deciding whether to proceed with the upgrade of the existing SLC3-compatible SLC4 worker nodes or wait for a new release.

   * We are suffering a lot because of two bugs in the LB service, which force us to restart the service from time to time (this week it happened twice). A brand new LB diagnostic was installed in PPS with Update26, so we offered to WMS developer to access our sever to debug. No reply came yet. 

   * FTS server was fixed and put back to production.   It is ready for testing FTS2.0 (currently in Certification)


  • 2007-04-25: Task: Problem with PPS SAM tests. Details: SAM and WMS ISes were(/are(?)) inconsistent. Due to this job submission for some sites (missing in ISM on WMS) was failing on WMS in workload manager at matchmaking with "BrokerHelper: Problems during rank evaluation (e.g. GRISes down, wrong JDL rank expression, etc.)". Problem debugged and understood. Status: Done (Follow up needed.) . Konstantin

  • 2007-04-24: Task: Many failure were observed for SAM replica management and SAM job submission tests on both lcg/glite CEs.
    Details: After testing the site manually no such failure was observed. Probably the cause of failure was due the use of production WMS and production SEs that were not published in PPS BDII.
    Status: None . Farida

  • 2007-04-24: Task: KickStart based PXE installation - problem! Details: Problem - short host name after host installation with "networking --bootproto=dhcp" in KickStart file. Problem confirmed for SLC308 and SLC 4X installers. Yvan will make a workaround in %post. The actual cause of the problem is not yet known!!! Status: Done . Konstantin

  • 2007-04-24: Task: VO testing SLC4 WN. Details: I gave interactive login to lxb1939 to users 'campanas' and 'rwalker' for Atlas. Status: Done . Antonio

  • 2007-04-24: Task: Add WMS lxb2070.cern.ch to WMS monitoring "http://lxb2007.cern.ch/pps_monitoring/pps_monitoring.html". Status: Done . Konstantin

  • 2007-04-24: Task: FTS lxb2064 (pps-fts.cern.ch) fixed. Details: However, tested between DPMs only. Added to monitoring in GOCDB. Ready to FTS2.0 update (now on CTB). Status: Done . Konstantin

  • 2007-04-24: Task: WMS testing. Details: bug #22337 (LB under-performing): I replied to the developer offering to provide data from our WMS LB for analysis. Status: Done . Antonio

  • 2007-04-23: Task: Host certificate expired. Details: lxb2006, new certificate in place Status: Done . Antonio

  • 2007-04-19: Task: PPS U27 - FTS, MON and E2EMONIT, WMS Status: Done . Konstantin

  • 2007-04-19: Task: problem with KickStart based installation: "network --bootproto=dhcp" in /afs/cern.ch/project/gd/yaim-server/kickstart/*.cfg leads to short host name assigned to a host. Checking. Status: Progress . Konstantin


CERN_PPS (C5 Report on 19-Apr-07)

CERN GRID PRE-PRODUCTION SITE (CERN_PPS) 
========================================

* PPS Update 27: done
     - BDII, LFC, VOBOX(Alice), lcg-RB, WMS and Torque batch updated 
     - test to upgrade from  SLC4(SLC3) to SLC4(SLC4) WNs done, The test failed 
        (the upgrade was not supported) and the results were sent to FIO for information

* Alice is testing natively compiled  SLC4 WNs on our internal batch system (torque)

* HDD failure on pps-wms (Workload Management Service). The node was retired and 
   replaced. A firewall problem on the new installation blocked job submission from 
   outside CERN for some hours. This had an impact on the submission of SAM tests 
   for the PPS
  

CERN_PPS (Section Report on 19-Apr-07)

   * PPS Update 27: in progress
      - BDII, WMS and Torque batch updated
      - test to upgrade from  SLC4(SLC3) to SLC4(SLC4) WNs done, The test failed 
        (the upgrade was not supported) and the results were sent to FIO for information

   * Alice is testing natively compiled  SLC4 WNs on our mini-batch system (torque)

   * FIO accepted to publish their pre-installation of natively compiled  SLC4 WNs in CERN_PPS.
      Schedule for the exercise to be delivered by the end of this week

   * A new WMS was installed (lxb2070.cern.ch) because of a HDD failure on lxb2092.cern.ch. 
      The dns alias pps-wms was changed accordingly. A firewall problem on the new installation 
      blocked the SAM submission for a while, now fixed.
  

  • 2007-04-19: Task: Firewall problem found in new WMS installation (lxb2070.cern.ch). Packets sent to Condor port (UDP) are blocked by the external firewall although configured to be open with the FW tool. Now fixed. Status: Done . Konstantin

  • 2007-04-18: Task: pps-bdii PPS-UPDATE 27. Status: Done . Lanxin Ma

  • 2007-04-18: Task:Alice test on gLite 3.1SLC4 . Status: Progress . Farida

  • 2007-04-18: Task: PPS-UPDATE 27. Status: Progress . Farida

  • 2007-04-18: Task: Some tests with gLite 3.1SLC4 batch nodes. Status: Done . Farida

  • 2007-04-18: Task: Test for upgrade from SL3 on SL4 WNs. Status: Done . Farida

  • 2007-04-18: Task: Installation of gLite 3.1 on SLC4 WNs from scratch. Status: Done . Farida

  • 2007-04-18: Task: Closed three tickets assigned to CERN_PPS. Status: Done . Farida

  • 2007-04-17: Task: new WMS installation (lxb2070.cern.ch). Details:HDD failed on lxb2092.cern.ch. Status: Done . Konstantin

  • 2006-09-20(5): Task: SAM in PPS. Details:SRM sensor installed Status: Done . Antonio

  • 2006-09-20(4): Task: SAM in PPS. Details:SE sensor installed Status: Done . Antonio

  • 2006-09-20(3): Task: SAM in PPS. Details:CE sensor installed Status: Done . Antonio

  • 2006-09-20(2): Task: SAM in PPS. Details:New user certificate requested to submit tests on behalf of OPS. Status: Done . Antonio

  • 2006-09-20(1): Task: SAM in PPS. Details:gCE sensor installed Status: Done . Antonio

  • 2006-09-19(1): Task: SAM in PPS. Details:Sam client installed Status: Done. Antonio

  • 2006-09-15(4): Good News : SFT tests in PPS now run on behalf of 'ops' vo

  • 2006-09-15(3): Task: SFT in PPS. Details:Migration to OPS VO - GIIS at PIC fixed. OPS VO active. Status: Done . Antonio

  • 2006-09-15(2): Task: SFT in PPS. Details:Migration to OPS VO - Request sent to PIC to support ops in their catalog, accepted by Carlos Status: Progress . Antonio

  • 2006-09-15(1): Task: SFT in PPS. Details:Migration to OPS VO - All SFT clients reconfigured. Status: Done . Antonio

  • 2006-09-14(2): Task: CERN PPS Site Admin. Details: kernel upgrade on lxb2093 Status: Done. Antonio.

  • 2006-09-14(1): Task: SFT in PPS. Details:Registration to Ops VO. Status: Done . Antonio

=> C5 report


  • 2006-08-28(1): Task: SFT in PPS. Details:Tests now run by UPATRAS. Status: Done . Antonio


=> C5 report

  • 2006-08-16(1): Task: CERN PPS Site Admin. Details: picard VO enabled on UI lxb1908 Status: Done. Antonio.

  • 2006-08-15(2): Task: SFT in PPS. Details:sft-apel test updated as requested by Frederich. Status: Done . Antonio

  • 2006-08-15(1): Task: CERN PPS Site Admin. Details: Script to monitor the ism set in crontab as for the Cern ROC WMS (lxb2059, pps-wms) Status: Done. Antonio.

  • 2006-08-11(3): Task: SFT in PPS. Details:Old client instance re-activated as back-up to replace temporarily main service at UPATRAS Status: Done. Antonio

  • 2006-08-11(2): Task: SFT in PPS. Details: GocSite_v04 and GocNode_v04 tables re-built after R_GMA registry upgrade at cesga. (again!) Status: Done. Antonio

  • 2006-08-11(1): Task: CERN PPS Site Admin. Details: Alias pps-wms requested to NETOPS for new CERN-PPS WMS. Status: Done. Antonio.


  • 2006-08-09(1): Task: PPS Coordination. Details: Reminder sent to PPS sites. the version will be forwarde on Monday, unless major issues are found Status: Done. Antonio.

  • 2006-08-08(2): Task: SFT in PPS. Details: CA test updated Status: Done. Antonio

  • 2006-08-08(1): Task: SFT in PPS. Details: GocSite_v04 and GocNode_v04 tables re-built after R_GMA registry upgrade at cesga. Status: Done. Antonio

  • 2006-08-04(1): Task: PPS Coordination. Details: Broadcast sent to open PPS version 3.0.2 to users. Status: Done. Antonio.

  • 2006-08-03(1): Task: PPS Coordination. Details: Mail sent to COD to re-start monitoring PPS after upgrade to gLite 3.0.2 . Status: Done. Antonio.

  • 2006-08-02(1): Task: CERN PPS Site Admin Details:Several Apel configuration problems fixed on CEs/ MONbox Status: Done . Antonio.

  • 2006-08-01(2): Task: PPS Coordination. Details: Mail sent to PPS to upgrade to gLite 3.0.2 Status: Done . Antonio.

  • 2006-08-01(1): Task: CERN PPS Site Admin Details: lxb2093 (MON) upgraded to gLite3.0.2. Status: Done . Antonio.

  • 2006-07-24(1): Task: FCR in PPS Details: Conf file generated by FCR is now copied in the central BDII location. Script set in my crontab. Status: Done . Antonio.

  • 2006-07-20(3): Task: SAM in PPS. Details: Oracle account EGEE_PPS_SAME requested Status: Done . Antonio

  • 2006-06-20(2): Task: SFT in PPS. Details: CA test upgraded (again) Status: Done . Antonio

  • 2006-06-20(1): Task: SFT in PPS. Details: CA test upgraded (altough no more active, the PPS instance can be useful for internal tests) Status: Done . Antonio.

  • 2006-06-15(1): Task: CERN PPS Site Admin. Details: Andrea's hack applied in the lxplus UI to copy central vomses file in local directory. Change request opened in Savannah and wiki page created in PPS wiki. Status: Done . Antonio.

  • 2006-05-31(3): Task: PPS Coordination. Details: PPS Service Status Spreadshhed upgraded (v18) and sent to sites for comments. Status: Done . Antonio.

  • 2006-05-31(2): Task: PPS Coordination. Details: New AFS UI broadcasted. Status: Done . Antonio.

  • 2006-05-31(1): Task: SFT in PPS Details: PPS acron jobs disabled. UPATRAS takes over in SFT maintanance. Status: Done . Antonio.

  • 2006-05-30(1): Task: CERN PPS Site Admin Details: Certificate of lxb1923 replaced Status: Done . Antonio.

  • 2006-05-29(1): Task: SFT in PPS Details: sft-lcg-rm-cr test fixed (in cvs) to solve the problem of contemporary submission from two CEs to single workernode. timestamp extended to seconds. Status: Done . Antonio.

  • 2006-05-26(2): Task: FCR in PPS Details: Broadcast sent to announce the FCR service for PPS. Status: Done . Antonio.

  • 2006-05-26(1): Task: SFT in PPS Details: Broadcast sent to announce that from Monday the SFT service will be run by UPATRAS. Status: Done . Antonio.

  • 2006-05-24(1): Task: Service configuration Details: configuration of owner-writer accounts for PPS FTS server/agents done with Konstantin Status: Done . Antonio.

  • 2006-05-23(1): Task: SFT in PPS Details: Central LFC at PIC went down --> SFT failed preliminary tests. SFT modified to use local catalog. Change made in cvs. Status: Done . Antonio.

  • 2006-05-22(3): Task: Documentation. Details: SFT on PPS wiki page changed to reflect the new merged SFT infrastructure. The new document has been passed to UPATRAS. Status: Done . Antonio.

  • 2006-05-22(1): Task: SFT in PPS. Details: SFT env test fixed. Now also the glite branch uses a real SE. Modification in CVS done by Frederic Schaer. Status: Done . Antonio.

  • 2006-05-19(1): Task: SFT in PPS. Details: SFT tests for gLite and LCG CEs merged to use gLite WMS Status: Done . Antonio.

  • 2006-05-18(1): Task: Documentation. Details: Site enrollment procedure moved to public wiki. Status: Done . Antonio.

  • 2006-05-17(2): Task: FCR in PPS. Details: Asked Judit for a quick implementation of a copy of the existing version of FCR. Status: Done . Antonio.

  • 2006-05-17(1): Task: SFT in PPS. Details: Certification Authorities test upgraded. Status: Done . Antonio.

  • 2006-05-11(1): Task: SFT in PPS. Details: UPATRAS agrees to run the SFT clients for pre-production. Status: Done . Antonio.

  • 2006-05-10(1): Task: SFT in PPS. Details: e-mail sent to George Goulas to ask UPATRAS to run the SFT clients for pre-production. Status: Done . Antonio.

  • 2006-04-25(1): Task: CERN PPS Site Admin Details: Mail sent to PPS users to open the service for RC2 Status: Done . Antonio.

  • 2006-04-20(3): Task: PPS Coordination. Details: Mail sent to PPS users to open the service for RC2 Status: Done . Antonio.

  • 2006-04-20(2): Task: SFT in PPS. Details: MONBOX upgraded. now uses new registry. SFT restored Status: Done . Antonio.

  • 2006-04-20(1): Task: PPS Coordination. Details: Central location defined for PPS BDII. A new afs group aretico:bdiiadm has been created to allow MAtteo Selmi and Daniele Cesini to maintain the BDII configuration file, tp be now kept under the directory /afs/cern.ch/project/gd/egee/www/preproduction/bdii Status: Done . Antonio.

  • 2006-03-24(1): Task: SFT in PPS. Details: cronjob added. Error message still sent to my inbox but seems not to affect the service. Installation documented on the wiki. Status: Done . Antonio.

  • 2006-03-22(1): Task: CERN PPS Site Admin Details: LFC catalog cunfiguration fixed. Now dilligent users can log in. Status: Done . Antonio

  • 2006-03-21(2): Task: Documentation Details: New page created in this wiki: How to installa and configure an AFS UI at CERN Status: Done . Antonio.

  • 2006-03-21(1): Task: CERN PPS Site Admin Details: AFS UI fixed. Now users can log in. Status: Done . Antonio

  • 2006-03-20(1): Task: SFT in PPS. Details: LCG instance now running. Status: Done . Antonio.

  • 2006-03-20(1): Task: PPS Coordination. Details: Phone call to Javier to discuss requirements for user info page Status: Done . Antonio.

  • 2006-03-17(1): Task: PPS Coordination. Details: Mail sent to experiments with Opening details for the PPS. Status: Done . Antonio.

  • 2006-03-14(2): Task: PPS Coordination. Details: Mail sent to Sven Gabriel with clarification of requirements for PPS sites. Status: Done . Antonio.

  • 2006-03-14(1): Task: PPS Coordination. Details: Spreadsheet with VO requirements changed with CMS requirements and circulated to Operations, Pre-production, SC, VOs Status: Done . Antonio.

  • 2006-03-13(2): Task: PPS Coordination. Details: Spreadsheet with VO requirements created and circulated to Operations, Pre-production, SC, VOs Status: Done . Antonio.

  • 2006-03-13(1): Task: PPS Coordination. Details: PPS operations meeting called and Agenda created Status: Done . Antonio.

  • 2006-03-07(2): Task: Documentation Details: New page created and pre-populated in the PPS wiki: Feedback on Middleware from sites Status: Done . Antonio.

  • 2006-03-07(1): Task: CERN PPS Site Admin Details: Several gLite3 services installed and configured. Several bugs opened. Status: Progress . Antonio. Alessandro, Constantin, Rolandas

  • 2006-03-06(2): Task: CERN PPS Site Admin Details: Several gLite3 services installed and configured. Several bugs opened. Status: Progress . Antonio. Alessandro, Constantin, Rolandas

  • 2006-03-06(1): Task: Ops Meeting. Details: Minutes taken Status: Done . Antonio.

  • 2006-03-03(3): Task: CERN PPS Site Admin Details: Several gLite3 services installed. Status: Progress . Antonio. Alessandro, Constantin, Rolandas

  • 2006-03-03(2): Task: PPS Coordination. Details: Status report for section meeting prepared Status: Progress . Antonio.

  • 2006-03-02(1): Task: PPS Coordination. Details: PPS service spreadsheet prepared Status: Progress . Antonio.

  • 2006-03-01(3): Task: SFT in PPS Details: first two sft tests job sent to CERN and LIP. Waiting for the display from Piotr. Status: Progress . Antonio.

  • 2006-03-01(2): Task: SFT in PPS Details: found an issue in new version of sftests: it suffers from a bug from lcg-infosites, addressed with gLite3 Workaround in place. Status: Progress . Antonio.

  • 2006-03-01(1): Task: SFT in PPS Details: LFC entries missing in CNAF BDII. Installation of a site BDII which should solve the problem agreed with PIC Status: Progress . Antonio.

  • 2006-02-23(1): Task: PPS Coordination. Details: e-mail sent to PPS calling for volonteers to install LCG-2 CEs Status: Done . Antonio.

  • 2006-02-22(4): Task: SFT in PPS Details: new version of the gocdb-xfer.py script with fixes for R-GMA installed Status: Done . Antonio.

  • 2006-02-22(3): Task: Documentation Details: procedure to install an LSF CE in PPS-CERN inserted in the wiki Status: Done . Antonio.

  • 2006-02-22(2): Task: Documentation Details: "Getting Started" procedure to work in PPS-CERN inserted in the wiki Status: Done . Antonio.

  • 2006-02-22(1): Task: PPS Coordination. Details: new mailing list for CERN_CIC created. Status: Done . Antonio.

  • 2006-02-21(3): Task: Accounting in PPS Details: Set-up of APEL schema started in CESGA. Status: Done . Antonio.

  • 2006-02-21(2): Task: SFT in PPS Details: LFC entries missing in CNAF BDII. Requested to Daniele Status: Progress . Antonio.

  • 2006-02-21(1): Task: SFT in PPS Details: APEL tables missing in PPS registry. Created by Javier Status: Progress . Antonio.

  • 2006-02-20(2): Task: SFT in PPS Details: LCG SFT instance:SQL problems introduced by R-GMA 1.5 worked around. Now the SFT query can be run. Flexible Archiver used; wiki page updated. Status: Progress . Antonio.

  • 2006-02-17(1): Task CERN PPS Site Admin Details: Kernel upgrade done to certification and preproduction clusters. Status: Done . Antonio.

  • 2006-02-16(1): Task GSTAT in PPS Details: E-mail sent to Min Tsai to obtain a PPS page in Gstat. Status: Done . Antonio.

  • 2006-02-15(2): Task SFT in PPS Details: Query in the SFT client for gLite changed, now it collects only certified sites. Status: Done . Antonio.

  • 2006-02-15(1): Task Bug Opened. Details: found a misbehaviour of voms-proxy-init. Bug #14997 submitted and assigned to Vincenzo Ciaschini Status: Done . Antonio.

  • 2006-02-09(3): Task SFT in PPS. Details: set-up of a separate SFt publisher started. Conflict found between MON and SFT rpms. Status: Progress . Antonio.

  • 2006-02-09(2): Task PPS Coordination. Details: e-mail sent to Pierre Girard (Lyon) to ask for upgrade FTS to 1.5 Status: Done . Antonio.

  • 2006-02-09(1): Task PPS Coordination. Details: e-mail sent to Mario David and Aveiro site (portugal) with a proposal for PPS support to Biomed. Currently being revised by LIP managers Status: Done . Antonio.

  • 2006-02-08(1): Meeting Christoph Witzig from SWITCH met. Successful integration of their CE with the PPS WMS at CERN. They should join the CERN ROC and create a site on the GOC dB to be monitored. Alessandro. Antonio.

  • 2006-02-07(1): Task PPS Coordination. Details: e-mail sent to Graeme Stewart (Glasgow) to ask for upgrade FTS to 1.5 Status: Done . Antonio.

  • 2006-01-26(5): Task PPS Coordination. Details: e-mail sent to Massimo Lamanna to regulate use of PPS WMSs by the ARDA monitor. Status: Done . Antonio.

  • 2006-01-26(5): Task CERN site troubleshooting Details: Debug of lxb2085 (WMS) failing the ARDA monitoring tests. Help from the developers to be required. Status: Progress . Antonio, Alessandro

  • 2006-01-26(4): Task CERN site Admin. Details: New failure in restarting of the R-GMA archiver after the power cut. The bug is still in "none" status. Entry in the PPS wiki page for this problem created. Status: Done . Antonio

  • 2006-01-26(3): Task CERN site Admin. Details: Old version of CA found on the R-GMA registry. For some reasons the old problem with the ca_eugridpma rpm was still there. The security utils rpm is still dependent on this. I needed to install the rpm from JRA1 repository to have the upgrade made. Status: Done . Antonio

  • 2006-01-26(2): Task CERN site Admin. Details: Yumit client configuration fixed on three nodes (LSF CE and R-GMA registry). Status: Done . Antonio

  • 2006-01-24(1): Task PPS Coordination. Details: Contact established with Yves Coppens (BHAM) willing to join the PPS. Email sent with pointer to the wiki page Status: Done . Antonio

  • 2006-01-23(1): Task PPS Coordination. Details: Preliminary plan for PPS evolution done and sent to PPS mailing list. Status: Done . Antonio

  • 2006-01-17(1): Task PPS troubleshooting Details: Help provided to FZK to fix R-GMA client configuration. Two bugs opened to gLite as a consequence. Status: Done . Alessandro, Sven Herrmann, Antonio

  • 2006-01-16(1): Task SFT in PPS. Details: lxplus062 taken out of production. gocdb-xfer cron job failing consequently. e-mail sent to Dave Kant to authorize new node. Status: Done . Antonio

  • 2006-01-11(1): Task Inter-section activity. Details: LCG2-Manual-Install changed for LCG2_7_0 Status: Done . Antonio

  • 2006-01-09(1): Task PPS site certification procedure. Details: e-mail sent to Rafal Lichwala to understand the current certification process. Status: Done . Antonio

  • 2006-01-06(1): Task CERN site Admin. Details: R-GMA server and LSF CE recovered after 01-05 power failure . Status: Done . Antonio

  • 2005-12-21(1): Task PPS Administration. Details: e-mail sent to Helene Cordier to organize PPS meeting in Barcelona. Status: Done . Antonio

  • 2005-12-19(1): Task CERN site Admin. Details: Host Certifcates expired on R-GMA registry lxb2087. Replaced (Thay had already been requested). Status: Done . Antonio

  • 2005-12-16(1): Task SFT in PPS. Details: SFT publisher was empty. Two main problems: 1)error in the cronjob for client; 2) authorization problems for server. Help to Judith to re-populate production R-GMA Status: Done . Antonio

  • 2005-12-15(1): Task PPS troubleshooting Details: Troubleshooting of lxb2090 (PPS CE): it now passes SFT Status: Done . Alessandro

  • 2005-12-14(2): Task PPS troubleshooting Details: Further troubleshooting of CNAF voms server (number of VOs is a problem) Status: Done . Alessandro

  • 2005-12-14(1): Task SFT in PPS. Details: account 'webman' on GOC db was deleted; as a consequence our instance of th goc-xfer.py script was failing. New account requested. The new account is enabled to connect only form lxplus062.cern.ch . acron job changed accordingly. Wiki page changed accordingly. Status: Done . Antonio

  • 2005-12-13(2): Task PPS troubleshooting Details: Troubleshooting of CNAF voms server (identified the problem) Status: Done . Alessandro

  • 2005-12-13(1): Task PPS troubleshooting Details: Troubleshooting of site Taiwan-PPS (job submission now works) Status: Done . Alessandro

  • 2005-12-12(3): Task PPS Coordination. Details: Mail sent to PPS to install QF "gLite QF1.4.1_31_2005" Status: Done . Antonio

  • 2005-12-12(2): Task PPS Coordination. Details: Minutes of 29-11-2005 PPS meeting copied in SA1 docs directory. Status: Done . Antonio

  • 2005-12-12(1): Meeting: Assessment with Hui-Min Lin about Taiwan-PPS status. Taiwan-PPS CE is now up and running. Hui-Min will check with the production managers of ASCG whether access can be given to production farms. A d-cache SE will be set-up to be tested with FTS. Antonio, Hui-Min Lin

  • 2005-12-9(1): Meeting: PPS and certification meeting to discuss the VOMS server issues. Maite Barroso Lopez; Maria Dimou; Maria Alandes Pradillo; Antonio Retico; Alessandro Usai; Nicholas Thackray; Di Qing; Joachim Flammer; Ian Neilson; Oliver Keeble

  • 2005-12-8(1): Task SFT in PPS. Details: goc-xfer.py script was failing because of the field 'grp' removed from the GOC db . As a consequence SFT tests were running on an empty list (not running). Script changed to new version. SFT run manually. Status: Done . Antonio

  • 2005-12-7(4): Task CERN site Admin. Details: Copy of the CERN PPS configuration files updated on the wiki page. Status: Done . Antonio

  • 2005-12-7(3): Good News : First "completely green" screen in the PPS SFT report page. By All PPS sites.

  • 2005-12-7(2): Task PPS Coordination. Details: e-mail sent to 'lcg-site-admin@hepNOSPAMPLEASE.ph.bham.ac.uk' to ask details about the client causing the WMS hanging problem. Ian Neilson and Romain Wartel in CC Status: Done . Antonio

  • 2005-12-7(1): Meeting: Assessment with Ian Neilson and Romain Wartel about the WMS hanging issue caused by external client. Nick, Antonio, Alessando, Ian, Romain.

  • 2005-12-6(3): Task CERN site Admin. Details: CERN WMS again up and running. Francesco Prelz found a client establishing continuous connections to the NS. Stifling the connection through the firewall made the trick. Now the matchmaking is fast enough Status: Done . Antonio

  • 2005-12-6(2): Task CERN site Admin. Details: CERN R_GMA registry and schema server added to the the goc db Status: Done . Antonio

  • 2005-12-5(1): Task - Manual Test. Details: PPS application to manage the sites out of the goc db Status: IN PROGRESS . Alessandro, Antonio

  • 2005-12-6(1): Task: PPS Coordination. Details: Mail sent to PPS to change the R-GMA registry to the CERN one. Status: Done. Antonio

  • 2005-12-5(2): Task: SFT in PPS. Details: PPS AFS UI at CERN reconfigured to use local registry, in order to re-activate SFT monitoring waiting for CESGA to come up again. Status: Done. Antonio

  • 2005-12-5(1): Task: SFT in PPS. Details: MON box at CERN reconfigured to run local registry and schema, in order to re-activate SFT monitoring waiting for CESGA to come up again. Status: Done. Antonio

  • 2005-12-2(2): Meeting: Assessment with David Collados about new i/o server SFT tests. No progress to report. Antonio. David Collados

  • 2005-12-2(1): Meeting: Kick-start Briefing with Hui-Min Lin from Taiwan-PPS. Antonio. Alessandro. Hui-Min Lin

  • 2005-12-1(7): Task: PPS Coordination. Details: (AP 2005.11.29-3) Overall check on the GOC db to find inconsistencies with the PPS wiki pages. Several Site Admins contacted via e-mail Status: Done. Antonio

  • 2005-12-1(6): Task: Documentation. Details: CERN Wiki updated: Now the CernNodeList is a link to GOC db Status: Done. Antonio

  • 2005-12-1(4): Task: SFT in PPS. Details: Cron Job changed: Now SFT tests are submitted every two hours. Status: Done. Antonio

  • 2005-12-1(3): Task: Documentation. Details: Link to new ScotGrid FTS service added in Core Services Status: Done. Antonio

  • 2005-12-1(2): Task: PPS Coordination. Details: Mail sent to UOA to get their wiki page updated. Status: Done. Antonio

  • 2005-12-1(1): Task: PPS Coordination. Details: Mail sent to Frederic Hemmer to have someone look at the CERN WMS, still slow. Status: Done. Antonio

  • 2005-11-30(5): Task: CERN site Admin. Details: Certificate renewal requested for three machines. Status: Done. Antonio

  • 2005-11-30(3): Task: Documentation. Details: Manual workaround obsoleted by quick fixes removed from the wiki page. Status: Done. Antonio

  • 2005-11-30(2): Task: CERN site Admin. Details: QF1.4.1_27_2005 and QF1.4.1_29_2005 installed on WNs and CEs. Regression test, bug closed and green light to PPS for installation sent. Status: Done . Antonio

  • 2005-11-30(1): Task: CERN site Admin. Details: Monbox configuration changed to patch the R-GMA Pong servlet security issue. Status: Done . Antonio

  • 2005-11-29(1): Meeting: PPS Operation Meeting (VRVS). PPS Site Admins

  • 2005-11-28(1): Task: SFT in PPS. Details: example grid-mapfile created for PPS to support VOMS server in version 1.3 and 1.4 . Communication sent to PPS mailing list Status: Done . Antonio

  • 2005-11-25(2): Task: SFT in PPS. Details: VOMS server of the SFT client changed to voms.cern.ch. Status: Done . Antonio

  • 2005-11-25(1): Task: SFT in PPS. Details: Query modified in the SFT client: now it checks the "Monitoring" flags. Status: Done . Antonio

  • 2005-11-24(3): Task: SFT in PPS. Details: Bug in the csh environment setenv found with the help of Maarteen and critical bug submitted to Savannah. Status: Done . Antonio

  • 2005-11-24(2): Task: PPS coordination. Details: Plan for the upgrade to 1.4.1 prepared and circulated to PPS. VRVS meeting booked. Status: Done . Antonio

  • 2005-11-24(1): Task: PPS coordination. Details: Agenda for next PPS Operations Meeting prepared (http://agenda.cern.ch/age?a057773) and circulated. VRVS meeting booked. Status: Done . Antonio

  • 2005-11-23(3): MEETING: PPS and certification meeting. Nick, Alessandro, Antonio, Di

  • 2005-11-23(2): Task: PPS Support Details: PPS troubleshooting with CESGA and CNAF Status: IN PROGRESS . Alessandro

  • 2005-11-23(1): Task: PPS Site Debugging. Details: WMS upgrade to 1.4.2 planning Status: IN PROGRESS . Alessandro

  • 2005-11-22(5): MEETING: PPS and certification meeting. Alessandro, Antonio, Daniele (CNAF)

  • 2005-11-22(4): Task: PPS Site Debugging. Details: grid-mapfile corrected on CERN CE. Status: Done . Antonio

  • 2005-11-22(3): Task: PPS CERN Site Administration. Details: CA Updated in CERN PPS CE and WNs. Status: Done . Antonio

  • 2005-11-22(1): Task: SFT in PPS. Details: Problems to job submission fixed. They were due to the host certificate of lcg-voms not copied on CNAF RB. It was causing a segmentation fault. Problem fixed and bug opened (#13734) Status: Done . Antonio

  • 2005-11-21(1): Task: Troubleshooting Details: PPS troubleshooting (apt repository at CNAF and CAs) Status: IN PROGRESS . Alessandro

  • 2005-11-18(1): Task: Troubleshooting Details: PPS troubleshooting Status: IN PROGRESS . Alessandro

  • 2005-11-18(2): Meeting: Integration of JRA1 tests in SFT. Details: existing IO and catalog tests will be integrated. David will get in phase with Javier. Assessment on Friday 2nd December. Status: Done . Antonio, Piotr, David Collados

  • 2005-11-18(1): Task: Documentation . Details: Procedure created on wiki to upgrade CA using LCG ones Status: Done . Antonio

  • 2005-11-17(3): Task: SFT in PPS. Details: SFT now runs twice per day in an acronjob (aretico) Status: Done . Antonio

  • 2005-11-17(2): Task: SFT in PPS. Details: Certificate with -noglobus option requested to CERN CA. Status: Done . Antonio

  • 2005-11-17(1): Task: SFT in PPS. Notes: Client intallation upgraded to fix RGMA test. Status: Done . Antonio, Piotr

  • 2005-11-04(2): Good News : SFT working in pre-production. By Antonio and Piotr

  • 2005-11-04(1): Task: SFT in PPS. Notes: Configured with Piotr's help Status: Done . Antonio, Piotr

  • 2005-11-02(1): Task: SFT in PPS. Notes: A bug was affecting the archiver in version 1.4. Fixed manually. Now sites and nodes are on the PPS R_GMA Status: Progress. Antonio

  • 2005-10-12(1): Task: SFT in PPS. Notes: The server now can be reached, but the data are stored in rgms as "latest". A new instance of the archiver needs to be created. Status: Progress. Antonio

  • 2005-10-6(1): Good News : PIC's PPS site gets access to the LCG's batch system. By Carlos Borrego Iglesias

  • 2005-10-2(1): Task: 006 - Dissemination. Notes: Information sent to Graeme Stewart from Glasgow wishing to join the PPS Status: Done. Antonio

  • 2005-09-21(7): Good News : mirror of the APT repository for the PPS created at CNAF by Matteo Selmi

  • 2005-09-21(6): Task: 005 - Documentation Notes: Procedure to handle the start of a VO in the PPS added to the wiki Status: Done. Antonio

  • 2005-09-21(5): Task : Debugging of the CERN CE (problems with the CESEBinding parameter not in the rgma server at cesga) STATUS:WORK IN PROGRESS. Di,Alessandro

  • 2005-09-21(4): Task : Gathering of the Preproduction information for ARDA (services, WNs, CPUs, upgrade status). Alessandro, Antonio

  • 2005-09-21(3): Task : 007 - Debugging Notes: Preproduction troubleshooting. Status: Done. Di,Alessandro

  • 2005-09-21(2): Task : 007 - Debugging Notes: debugging of the SC at PIC and troubleshooting. Status: Done. Di,Alessandro

  • 2005-09-21(1): Task : 007 - Debugging Notes: Debugging of SFT and a new critical bug submitted as a consequence. Status: Done. Antonio, Piotr

  • 2005-09-20(2): Task : 007 - Debugging Notes: Debugging of the preproduction Io server. Status: Done. Di,Alessandro

  • 2005-09-20(1): Task: 004 - Critical bug report Notes: #10872 VO Software TAG publishing Status: Done. Di

  • 2005-09-19(4): Task : Debugging of the preproduction WMS (Maarten,Di,Alessandro,Antonio)

  • 2005-09-19(3): Good News : First CMS Analysis job accessing CMS data from Castor run at CERN by Stefano Belforte (CMS)

  • 2005-09-19(2): Meeting - CMS meeting to define the requirements for cms to run in the glite preproduction area. Alessandro, Antonio, Di, Stefano Belforte, Marco Verlato (phone)

  • 2005-09-19(1): Task: 003 - Apt repository Notes: Mirror created for hte pre-production area (CNAF) Status: Done. Alessandro

  • 2005-09-14(1): Meeting: Preproduction and CMS - Discussion about the possibility for cms to run on the preproduction sites. Present Nick, Di, Antonio and Alessandro.

  • 2005-08-29(2): Task: 002 - Send lsf_install.sh to developers for feedback Notes: Sent via e-mail to Frederic Hemmer. Status: Done. Antonio

  • 2005-08-29(1): Task: 001 - WMS not working: Analize WMS crash Notes: Malfunctioning seems not to be related to the same one at CNAF. The problem at CNAF was probably due to a connected CE (Daniele should report it as a bug) Status: Done. Antonio

  • 2005-08-24(2): Started installing FTS/FPS. Di

  • 2005-08-24(1): Fixed the problems of LSF CE. Di

  • 2005-08-19: Pre Production wiki page with work log started. Antonio

-- Main.dqing - 24 Aug 2005

Edit | Attach | Watch | Print version | History: r355 < r354 < r353 < r352 < r351 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r355 - 2011-06-21 - AndresAeschlimann
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback